# Robust Superradiance and Spontaneous Spin Ordering in Disordered Waveguide Quantum Electrodynamics

### v2

@@ I Introduction
Collective spontaneous emission, or superradiance, is a hallmark of cooperative quantum behavior, first predicted by Dicke in 1954 [1]. It was shown that the collective emission of a set of $N$ two-level atoms (TLAs) or other emitters can be drastically accelerated by the build-up of correlations between individual dipoles. In Dicke’s idealized model, with symmetric all-to-all interactions, this effect leads to a sharp superradiant photon burst at a time $\sim\ln(N)/(N\gamma)$, with a peak emission rate proportional to $\gamma N^{2}$, where $\gamma$ is the decay rate of an individual emitter. Since then, this phenomenon has sparked extensive theoretical [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] and experimental [12, 13, 14, 15] investigations, and remains a cornerstone of quantum optics.
Recently, there has been a revived interest in superradiance [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], mainly stimulated by the development of new experimental platforms, in which superradiant effects can be explored at an unprecedented level of control, and potentially be exploited in applications such as superradiant lasers [43, 44, 45] or quantum metrology [46, 47, 48, 49]. Of particular relevance in this context are waveguide QED systems [50, 51, 52], where the emitted photons are confined to a one-dimensional (1D) channel and atom-atom correlations can be established efficiently and over long distances [16, 18, 19, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Specifically, when emitters are placed in a lattice with a spacing that matches an exact multiple of the resonant wavelength, all the emitted photons can interfere fully constructively, thereby recovering the features of the original Dicke model in an extended setting. However, in real systems, the positions of the atoms as well as their frequencies fluctuate, giving rise to competing interference patterns once the disorder becomes sufficiently strong. This raises an important fundamental question: Does Dicke superradiance survive in disordered quantum systems, and how does disorder affect its characteristic features?
@@
While it is generally expected that collective radiation effects are relatively robust against weak disorder, the fate of superradiant emission under strong disorder remains unclear. This concerns, in particular, the scaling behavior of the superradiant burst in 1D, where superradiance could be rather robust due to long-range interactions on the one hand, while on the other hand, the influence of competing interference patterns or disorder-induced localization is most pronounced. In prior works, the existence of superradiance in disordered systems has been inferred from initial photon correlations [21, 28], upper bounds on the maximal decay rate [39] and heuristic scaling arguments [40], but the validity of these arguments and the actual scaling of the peak emission rate remain poorly understood. In particular, it is unclear whether the $N^{2}$ scaling survives in the presence of strong disorder, and if so, how the atoms can sustain collective behavior despite competing phase relations that disrupt spatial correlations and a global phase alignment.
@@ II Model
In our study, the offset phases $\xi_{j}=k_{0}\delta_{j}$ are randomly chosen from a uniform distribution in the interval $\Theta/2,\Theta/2)$, such that $\Theta$ quantifies the degree of disorder. Note that for this configuration, neighboring atoms can overlap when $\Theta>2\pi$ and in the limit $\Theta\rightarrow\infty$ we recover a fully homogeneous density distribution spatially. The dynamics is determined by the distribution of the relative propagation phases, $\Delta\xi_{ij}=(\xi_{i}-\xi_{j})\mod 2\pi$, which obeys a wrapped triangular distribution (see the Supplemental Material [60]) and approaches a uniform distribution for any $\Theta=2n\pi$ with $n\in\mathbb{Z}_{+}$ and in the limit $\Theta\rightarrow\infty$. Additional simulations for an ensemble with Gaussian-distributed $\xi_{j}$ are performed for comparison, ensuring that the general conclusions do not depend on any specifics of the disorder.
@@ III.2 Quantum State Diffusion
In general, the wavefunction in Eq. (III.2) still scales exponentially with the atom number, and to make this method tractable, we approximate each trajectory by a product state, $|\psi\rangle\approx\otimes_{j=1}^{N}|\psi_{j}\rangle$. This simplification allows us to describe each trajectory again in terms of classical spin variables $\vec{\mathfrak{s}}_{j}(t)$, with components $\mathfrak{s}^{\eta}_{j}=\langle\psi_{j}|\sigma_{j}^{\eta}|\psi_{j}\rangle$. The corresponding stochastic equations for these spin vectors for the current setting are given in Ref. [36]. Note that QSDMF can be improved by refining the treatment of individual trajectories. For example, including second-order moments [62, 63] in the EOM of trajectories can encode some quantum correlations, at the cost of a less favorable scaling $\mathcal{O}(N^{2})$. Importantly, by treating trajectories using matrix product states [36], it can be systematically improved by increasing the bond dimension $d$, with complexity $\mathcal{O}(Nd^{3})$. This method is thus particularly efficient if entanglement in each trajectory is constant or logarithmic in $N$, which requires $d\sim\mathcal{O}(1)$ or $d\sim\mathcal{O}(N)$, respectively, for giving numerically exact results.
@@ III.3 Benchmarking
The comparison shows excellent agreement between semiclassical simulations and quantum jump results, even for modest system sizes, as well as a systematic reduction of the relative errors for increasing $N$. This concerns, in particular, the location and height of the superradiant peak, which are the focus of our discussion below. For DTWA, the comparison becomes less accurate at the final stage of the decay, which can be attributed to the MF-like factorization of the photon-atom coupling in Eq. (9) and Eq. (10). However, we conclude that for the transient regime of interest, both the DTWA and the QSDMF simulations capture very accurately the collective decay dynamics, while being at the same time numerically highly efficient and scalable to hundreds or thousands of atoms.
@@ III.4 Comparison with Other Methods
To conclude, out of these various methods, the DTWA approach described above combines an efficient scaling, $\mathcal{O}(N)$, with the ability to account for both classical and—up to a certain degree—quantum correlations and permits a straightforward and numerically inexpensive generalization to non-Markovian settings. This makes this method suitable for studying superradiance in a large variety of different scenarios, and has also been selected as the method of choice for most parts of the numerical analysis below. Additionally, we use QSDMF to gain trajectory-level insights, for example, on the spontaneous spin ordering studied in Sec. V.
@@ IV.2 Finite-Size Scaling
While this analysis does not allow us to predict the numerical prefactor $r_{0}$, it shows that the observed scaling relations for the maximal decay rate during the burst can be obtained from the relation $\mathcal{R}_{\star}\propto N\Gamma_{+}$. This observation is rather surprising, as in view of the nonlinear dynamics or the presence of additional Hamiltonian terms, such a simple correspondence is a priori not expected. As discussed in Appendix E, this observation is, however, consistent with an upper bound that is proportional to $\Gamma_{+}N$.
Note that for any non-vanishing $\mu_{a}>0$, the anomalous finite-size scaling $\sim 1/\sqrt{N}$ is observed up to an atom number of $N\lesssim(\sigma_{a}^{2}+\sigma_{b}^{2})/\mu_{a}^{2}$, before there is a transition to a $1/N$ scaling for very large $N$. This explains the rather slow convergence for all disorder strengths $\Theta\gg 1$. In Appendix D, we also present $\mathcal{R}_{\star}$ for an ensemble of atoms with Gaussian-distributed phases $\xi_{j}$. Again, we find a disorder-dependent scaling toward the asymptotic limit, which can be explained by this relation between the mean value and the fluctuations in the propagation phases.
@@ V.1 Maximal Decay Rate
While superradiance was originally described in terms of transitions between highly entangled collective states [1], most features of this phenomenon can be explained in terms of classical correlations between the atomic dipoles. More specifically, it has been shown [36] that in general, when $N$ is large, the superradiant decay can be accurately described by a mixed state of the form,
@@
Recently, by applying a theorem proved by Lieb [67] and Bravyi et al. [68] for bounding the ground state energy of spin-$1/2$ Hamiltonians, an upper bound for $\mathcal{R}_{\star}$ has been derived in Ref. [39]. By refining this calculation, and after taking the disorder average, we obtain (see Appendix E)
@@
Note that the upper bound $\mathbb{R}$ is independent of the Hamiltonian $H$, which indicates that details in $H$ do not change the superradiant scaling. For example, as shown in Appendix F, artificially ignoring the Hamiltonian only modifies the prefactor of the $N^{2}$-scaling.
@@ V.2 Spontaneous Spin Ordering
Indeed, we find that both trends are clearly visible in the examples shown in Fig. 8, although considerably broadened by fluctuations. This confirms that, under strong disorder, the spins self-organize their orientations along a local direction determined by their random position. This is in stark contrast to the weakly disordered cases, where spins tend to synchronize with a common phase. Interestingly, under all conditions, we find a particularly strong alignment tendency for atoms that are multiples of a wavelength apart, i.e., $\Delta\xi=0$. In this case, there is an almost vanishing probability to find a nonzero relative phase $\Delta\phi$, between the corresponding dipoles. Note that these findings are not constrained to systems with disorder. In Appendix G, we perform an identical analysis for periodic arrays, where for lattice spacing that do not match the emission wavelength $\lambda_{0}$, similar orderings of the atomic dipoles arise.
@@ V.3 Mirror-Asymmetric Correlations
In Fig. 9, we plot the actual distribution of the pair of rates $(\mathcal{R}_{\text{R}},\mathcal{R}_{\text{L}})$ in each QSDMF trajectory at time $t=t_{\star}$. These plots provide a very intuitive way to illustrate the effect of mirror-asymmetric emission in superradiant decay. For weak disorder both rates are strongly correlated around $\mathcal{R}_{\text{R}}\approx\mathcal{R}_{\text{L}}$. This implies that, along each trajectory, photons are emitted approximately equally likely to the left and to the right. In the regime of strong disorder, the rates are anti-correlated with $\mathcal{R}_{\text{R}}+\mathcal{R}_{\text{L}}\approx\mathcal{R}_{\star}$. Therefore, in this regime, the atomic dipoles evolve into configurations that successively lead to preferred emission into either the right or the left. Additional simulations presented in Appendix H show that this preference exists during most parts of the decay, although individual trajectories may also change their preferred emission direction during the process.
In summary, Fig. 8 and Fig. 9 clearly highlight the difference in the spontaneously established ordering of the atomic dipoles in weakly and strongly disordered arrays. For strong disorder, the peaks in the diagonals of the distribution of $p(\Delta\xi,\Delta\phi)$ in Fig. 8(c) and the anticorrelations seen in the rates in Fig. 9(b) also demonstrate the relevance of those dipole orderings that maximize emission to the left or the right. However, as evident from the photon correlations and the distribution of spin orientations, these orderings are distributions with a considerable broadening around the optimal configurations, which is also the reason underlying the discrepancy between the upper bound $\mathbb{R}_{<}$ and actual numerical data shown in Fig. 7. Note that our analysis has focused on spin ordering in the bulk of the atomic array. In a related recent analysis of spin-ordering in pumped atomic arrays [69], it was observed that atoms at the left (right) end align anticlockwise (clockwise) and therefore prefer emission to the left (right). In our setup, such boundary effects also exist, but their contribution to the overall dynamics and decay asymmetry is minor (see Appendix I).

### v3

@@ I Introduction
Collective spontaneous emission, or superradiance, is a hallmark of cooperative quantum behavior, first predicted by Dicke in 1954 [1]. It was shown that the collective emission of a set of $N$ two-level atoms (TLAs) or other emitters can be drastically accelerated by the build-up of correlations between individual dipoles. In Dicke’s idealized model, with symmetric all-to-all interactions, this effect leads to a sharp superradiant photon burst at a time $\sim\ln(N)/(N\gamma)$, with a peak emission rate proportional to $\gamma N^{2}$, where $\gamma$ is the decay rate of an individual emitter. Since then, this phenomenon has sparked extensive theoretical [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and experimental [18, 19, 20, 21] investigations, and remains a cornerstone of quantum optics.
Recently, there has been a revived interest in superradiance [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], mainly stimulated by the development of new experimental platforms, in which superradiant effects can be explored at an unprecedented level of control, and potentially be exploited in applications such as superradiant lasers [50, 51, 52] or quantum metrology [53, 54, 55, 56, 57]. Of particular relevance in this context are waveguide QED systems [58, 59, 60], where the emitted photons are confined to a one-dimensional (1D) channel and atom-atom correlations can be established efficiently and over long distances [22, 24, 25, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. Specifically, when emitters are placed in a lattice with a spacing that matches an exact multiple of the resonant wavelength, all the emitted photons can interfere fully constructively, thereby recovering the features of the original Dicke model in an extended setting. However, in real systems, the positions of the atoms as well as their frequencies fluctuate, giving rise to competing interference patterns once the disorder becomes sufficiently strong. This raises an important fundamental question: Does Dicke superradiance survive in disordered quantum systems, and how does disorder affect its characteristic features?
@@
While it is generally expected that collective radiation effects are relatively robust against weak disorder, the fate of superradiant emission under strong disorder remains unclear. This concerns, in particular, the scaling behavior of the superradiant burst in 1D, where superradiance could be rather robust due to long-range interactions on the one hand, while on the other hand, the influence of competing interference patterns or disorder-induced localization is most pronounced. In the studies of pencil-shaped atomic gases [17, 18, 12, 13, 14, 15, 16, 11], which can be approximated as a 1D model, superradiance has been shown with the atoms treated as a continuous medium and without quantum fluctuations in the decaying dynamics. It remains unclear if superradiance can exist in a fully quantum model and beyond the continuum limit. Recently, the existence of superradiance in disordered systems has been inferred from initial photon correlations [28, 35], upper bounds on the maximal decay rate [46], and heuristic scaling arguments [47], but the validity of these arguments and the actual scaling of the peak emission rate remain poorly understood. In particular, it is unclear whether the $N^{2}$ scaling survives in the presence of strong disorder, and if so, how the atoms can sustain collective behavior despite competing phase relations that disrupt spatial correlations and a global phase alignment.
@@ II Model
In our study, the offset phases $\xi_{j}=k_{0}\delta_{j}$ are randomly chosen from a uniform distribution in the interval $\Theta/2,\Theta/2)$, such that $\Theta$ quantifies the degree of disorder. Note that for this configuration, neighboring atoms can overlap when $\Theta>2\pi$ and in the limit $\Theta\rightarrow\infty$ we recover a fully homogeneous density distribution spatially. The dynamics is determined by the distribution of the relative propagation phases, $\Delta\xi_{ij}=(\xi_{i}-\xi_{j})\mod 2\pi$, which obeys a wrapped triangular distribution (see the Supplemental Material [68]) and approaches a uniform distribution for any $\Theta=2n\pi$ with $n\in\mathbb{Z}_{+}$ and in the limit $\Theta\rightarrow\infty$. Additional simulations for an ensemble with Gaussian-distributed $\xi_{j}$ are performed for comparison, ensuring that the general conclusions do not depend on any specifics of the disorder. To connect the results to realistic platforms, in [68], we also discusses the expected range of the disorder strength $\Theta$ for a representative set of experimental systems with atomic and solid-state emitters.
@@ III.2 Quantum State Diffusion
In general, the wavefunction in Eq. (III.2) still scales exponentially with the atom number, and to make this method tractable, we approximate each trajectory by a product state, $|\psi\rangle\approx\otimes_{j=1}^{N}|\psi_{j}\rangle$. This simplification allows us to describe each trajectory again in terms of classical spin variables $\vec{\mathfrak{s}}_{j}(t)$, with components $\mathfrak{s}^{\eta}_{j}=\langle\psi_{j}|\sigma_{j}^{\eta}|\psi_{j}\rangle$. The corresponding stochastic equations for these spin vectors for the current setting are given in Ref. [43]. Note that QSDMF can be improved by refining the treatment of individual trajectories. For example, including second-order moments [70, 71] in the EOM of trajectories can encode some quantum correlations, at the cost of a less favorable scaling $\mathcal{O}(N^{2})$. Importantly, by treating trajectories using matrix product states (MPS) [43], it can be systematically improved by increasing the bond dimension $d$, with complexity $\mathcal{O}(Nd^{3})$. This method is thus particularly efficient if entanglement in each trajectory is constant or logarithmic in $N$, which requires $d\sim\mathcal{O}(1)$ or $d\sim\mathcal{O}(N)$, respectively, for giving numerically exact results.
@@ III.3 Benchmarking
The comparison shows excellent agreement between semiclassical simulations and quantum jump results, even for modest system sizes, as well as a systematic reduction of the relative errors for increasing $N$. This concerns, in particular, the location and height of the superradiant peak, which are the focus of our discussion below. For DTWA, the comparison becomes less accurate at the final stage of the decay, which can be attributed to the MF-like factorization of the photon-atom coupling in Eq. (9) and Eq. (10). However, we conclude that for the transient regime of interest, both the DTWA and the QSDMF simulations capture the collective decay dynamics very accurately, while being at the same time numerically highly efficient and scalable to hundreds or thousands of atoms. In addition, similar to Ref. [43], an argument for the underlying reason behind the accuracy of QSDMF for large $N$ is presented in Appendix D.
@@ III.4 Comparison with Other Methods
To conclude, out of these various methods, the DTWA approach described above combines an efficient scaling  [^2] Note that, for ease of comparison, the scalings here are for short-range interactions. For long-range interactions, the scalings are modified by additional $N$-dependent factors. In the all-to-all case considered here, DTWA, QSDMF, and CE each pick up an extra factor of $N$, whereas QSD+MPS picks up an extra factor of $N^{2}$., $\mathcal{O}(N)$, with the ability to account for both classical and—up to a certain degree—quantum correlations and permits a straightforward and numerically inexpensive generalization to non-Markovian settings. This makes this method suitable for studying superradiance in a large variety of different scenarios, and has also been selected as the method of choice for most parts of the numerical analysis below. Additionally, we use QSDMF to gain trajectory-level insights, for example, on the spontaneous spin ordering studied in Sec. V.
@@ IV.2 Finite-Size Scaling
While this analysis does not allow us to predict the numerical prefactor $r_{0}$, it shows that the observed scaling relations for the maximal decay rate during the burst can be obtained from the relation $\mathcal{R}_{\star}\propto N\Gamma_{+}$. This observation is rather surprising, as in view of the nonlinear dynamics or the presence of additional Hamiltonian terms, such a simple correspondence is a priori not expected. As discussed in Appendix F, this observation is, however, consistent with an upper bound that is proportional to $\Gamma_{+}N$.
Note that for any non-vanishing $\mu_{a}>0$, the anomalous finite-size scaling $\sim 1/\sqrt{N}$ is observed up to an atom number of $N\lesssim(\sigma_{a}^{2}+\sigma_{b}^{2})/\mu_{a}^{2}$, before there is a transition to a $1/N$ scaling for very large $N$. This explains the rather slow convergence for all disorder strengths $\Theta\gg 1$. In Appendix E, we also present $\mathcal{R}_{\star}$ for an ensemble of atoms with Gaussian-distributed phases $\xi_{j}$. Again, we find a disorder-dependent scaling toward the asymptotic limit, which can be explained by this relation between the mean value and the fluctuations in the propagation phases.
@@ V.1 Maximal Decay Rate
While superradiance was originally described in terms of transitions between highly entangled collective states [1], most features of this phenomenon can be explained in terms of classical correlations between the atomic dipoles. More specifically, it has been shown [43] that in general, when $N$ is large, the superradiant decay can be accurately described by a mixed state of the form (see also Appendix D),
@@
Recently, by applying a theorem proved by Lieb [76] and Bravyi et al. [77] for bounding the ground state energy of spin-$1/2$ Hamiltonians, an upper bound for $\mathcal{R}_{\star}$ has been derived in Ref. [46]. By refining this calculation, and after taking the disorder average, we obtain (see Appendix F)
@@
Note that the upper bound $\mathbb{R}$ is independent of the Hamiltonian $H$, which indicates that details in $H$ do not change the superradiant scaling. For example, as shown in Appendix G, artificially ignoring the Hamiltonian only modifies the prefactor of the $N^{2}$-scaling.
@@ V.2 Spontaneous Spin Ordering
Indeed, we find that both trends are clearly visible in the examples shown in Fig. 8, although considerably broadened by fluctuations. This confirms that, under strong disorder, the spins self-organize their orientations along a local direction determined by their random position. This is in stark contrast to the weakly disordered cases, where spins tend to synchronize with a common phase. Interestingly, under all conditions, we find a particularly strong alignment tendency for atoms that are multiples of a wavelength apart, i.e., $\Delta\xi=0$. In this case, there is an almost vanishing probability to find a nonzero relative phase $\Delta\phi$, between the corresponding dipoles. Note that these findings are not constrained to systems with disorder. In Appendix H, we perform an identical analysis for periodic arrays, where for lattice spacing that do not match the emission wavelength $\lambda_{0}$, similar orderings of the atomic dipoles arise.
@@ V.3 Mirror-Asymmetric Correlations
In Fig. 9, we plot the actual distribution of the pair of rates $(\mathcal{R}_{\text{R}},\mathcal{R}_{\text{L}})$ in each QSDMF trajectory at time $t=t_{\star}$. These plots provide a very intuitive way to illustrate the effect of mirror-asymmetric emission in superradiant decay. For weak disorder both rates are strongly correlated around $\mathcal{R}_{\text{R}}\approx\mathcal{R}_{\text{L}}$. This implies that, along each trajectory, photons are emitted approximately equally likely to the left and to the right. In the regime of strong disorder, the rates are anti-correlated with $\mathcal{R}_{\text{R}}+\mathcal{R}_{\text{L}}\approx\mathcal{R}_{\star}$. Therefore, in this regime, the atomic dipoles evolve into configurations that successively lead to preferred emission into either the right or the left. Additional simulations presented in Appendix I show that this preference exists during most parts of the decay, although individual trajectories may also change their preferred emission direction during the process.
The anti-correlation between $\mathcal{R}_{\text{R}}$ and $\mathcal{R}_{\text{L}}$ stands in stark contrast to the theoretical predictions at the MF level without dynamical quantum fluctuations [17, 13]. There, with fluctuations only in the initial condition, the left- and right-going emissions are predicted to be essentially symmetric and uncorrelated. This highlights the crucial role of the quantum fluctuations in our simulation methods, both in enabling accurate predictions and in revealing the underlying microscopic mechanism.
In summary, Fig. 8 and Fig. 9 clearly highlight the difference in the spontaneously established ordering of the atomic dipoles in weakly and strongly disordered arrays. For strong disorder, the peaks in the diagonals of the distribution of $p(\Delta\xi,\Delta\phi)$ in Fig. 8(c) and the anticorrelations seen in the rates in Fig. 9(b) also demonstrate the relevance of those dipole orderings that maximize emission to the left or the right. However, as evident from the photon correlations and the distribution of spin orientations, these orderings are distributions with a considerable broadening around the optimal configurations, which is also the reason underlying the discrepancy between the upper bound $\mathbb{R}_{<}$ and actual numerical data shown in Fig. 7. Note that our analysis has focused on spin ordering in the bulk of the atomic array. In a related recent analysis of spin-ordering in pumped atomic arrays [78], it was observed that atoms at the left (right) end align anticlockwise (clockwise) and therefore prefer emission to the left (right). In our setup, such boundary effects also exist, but their contribution to the overall dynamics and decay asymmetry is minor (see Appendix J).
@@ VII.1 Summary
VII.2 Relation to Previous Work
Since Dicke’s pioneering work [1], superradiance has been studied extensively in atomic gases and other extended media [17, 18, 12, 13, 14, 15, 16, 11]. In the continuum limit, where a dense ensemble can be approximated as a continuous medium rather than as discrete radiating dipoles, collective bursts are well established. However, much less is understood about superradiance beyond this limit in strongly disordered discrete systems, where permutation symmetry is absent and the many-body dynamics rapidly becomes intractable at large $N$. While recent studies [28, 35, 46] have provided interesting arguments about collective decay in extended or disordered arrays based on photon correlations or upper bounds of maximal decay rates, there is no direct large-$N$ characterization, let alone a microscopic mechanism for its robustness.
Building on the recent observation that local observables in superradiant decay can admit an accurate separable description [43], we address this problem using two complementary large-scale simulation methods together with an analytical approach. Beyond establishing the robustness of the superradiant $N^{2}$ scaling in the disordered waveguide setting, our central conceptual advance is to identify the mechanism that makes this robustness possible: a spontaneous disorder-aligned spin ordering generated by the competition between the left- and right-propagating decay channels. Relative to previous work, the present results contribute in two main ways.
(i) We provide direct large-$N$ evidence that the characteristic superradiant $N^{2}$ scaling remains intact even under strong spatial disorder in a bidirectional waveguide. At the same time, the trajectory-based simulations reveal fluctuation-induced effects that are absent in continuum or MF descriptions [17, 18, 12, 13, 14, 15, 16, 11], most notably the spontaneous selection of one of two ordered spin configurations and the resulting anti-correlation between left- and right-going emission channels. These results show that fluctuations are not merely quantitative corrections here, but are essential for capturing the microscopic physics of the collective burst.
(ii) To our knowledge, the spontaneous spin ordering identified here is a new self-organization mechanism for superradiant decay in waveguide QED, especially in the strongly disordered regime considered here. For zero or weak spatial disorder, phase synchronization into a nearly uniform configuration is well known in spontaneous emission [17] and under incoherent pumping [23]. Likewise, for purely chiral emission with a single decay channel, the propagation phases can be absorbed into a redefinition of spin operators and therefore do not affect the dynamics (see e.g. [4, 17, 37]). In contrast, in the bidirectional case studied here, the competition between two decay channels gives rise to two emergent spin configurations whose phases adapt to the random atomic locations and thereby preserve constructive interference. We further show that related spin ordering also appears in regularly spaced arrays, suggesting that this mechanism is not restricted to disordered ensembles and may arise more broadly in cooperative light-matter dynamics.
VII.3 Outlook